sir if x^3-31x+30=0 then how we will find values of x

sir


if x^3-31x+30=0


then how we will find values of x 

Grade:12

1 Answers

bibhash jha
13 years ago

As we can see clearly , x=1 is a solution . therefore (x-1) is a factor

so dividing x^3-31x+30 by (x-1) we get x^2+x-30

so , x^3-31x+30=(x-1)(x^2+x-30)= (x-1)(x+6)(x-5)

therefore roots are x=1,-6 and 5